OpenStudy (anonymous):

x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (compassionate):

What do you think you'll do first, Asc? I'll give you hint, use the distribution property.

OpenStudy (compassionate):

I'm willing to help you understand this, but if you're not willing to work with me then I won't solve it.

OpenStudy (anonymous):

x=-9 & x=-8?

OpenStudy (compassionate):

x can only be equal to one thing, let me do the first step for you. You see those brackets? Just multiply the number on the outside by the numbers on the inside. -3x -9 = -3x -3 -5 Now what do you think needs to be done?

OpenStudy (anonymous):

x = 1?

OpenStudy (compassionate):

x = 9 -3 -5 x = 1 good job! :)